OpenStudy (anonymous):

solve. 7x-17/8 = 4 / = fraction br

OpenStudy (anonymous):

7x - 17/8 = 4 ***multiply both sides by 8*** 56x - 17 = 32 56x = 32 + 17 56x = 49 x = 49/56

OpenStudy (anonymous):

i thought the answer is x=89/126 without multiplying 8 on both sides

OpenStudy (anonymous):

noope it should give the same answer 7x = 4 + 17/8 7x = 49/8 x = 49/56